# Positive-Unlabeled Constraint Learning (PUCL) for Inferring Nonlinear Continuous Constraints Functions from Expert Demonstrations

### v1

@@ IV-C Results of Different Policy Representation and Learning Methods
As introduced in section III-D, we considered two policy representing and learning approaches, namely CRL (Constrained Reinforcement Learning) and DSM (Dynamical System Modulation). We test the two variants in the 3D task and summarize the metrics at convergence and training time of CRL-PUCL and DSM-PUCL in Table II. Here, the recall rate is computed as (True infeasible)/(True infeasible + False feasible), and precision is (True infeasible)/(True infeasible + False infeasible). Training was conducted on a PC with a 12th Intel i9-12900K × 24 and an NVIDIA GeForce RTX 3070. All results are the average of 15 independent runs.
@@
DSM-PUCL and CRL-PUCL achieve similar classification performance in IoU, recall rate and precision, and DSM-PUCL generally exhibits lower variance. For the learned policy, while both methods have an unsafe rate below $1\%$, the rate of DSM-PUCL is much lower than CRL-PUCL. This is because DSM-PUCL modulates the policy in a principled way and features a safety guarantee [25]. DSM-PUCL also requires less training time than CRL-PUCL. Based on these comparison, we conclude that both methods are effective for learning constraints. For robotic tasks where the control variable is the velocity/acceleration of each state dimension, DSM may be more favourable than CRL in terms of training time and policy performance, while CRL is more suitable for more general task with more complex environment dynamics and rewards.
TABLE II: Comparison of two policy representing and learning methods
@@
IV-D Constraint Transfer to Variant of the Same Task
One remarkable advantage of learning constraint is that the learned constraint can be transferred to similar task with the same constraint but potentially different goals and rewards. To demonstrate this, we consider the 3D tasks of robot avoiding four tall cylinders forming a convex shape (see Fig. 1). We first applied DSM-PUCL to learn this constraint from 19 demonstrated trajectories (Fig. 1 shows some of the demonstrations.) Then the learned constraint network is transferred to a relevant task with a shifted goal state $s_{g}^{\prime}$ and a set of different starting points. The new nominal policy heading to $s_{g}^{\prime}$ is directly modulated with the transferred constraint, still following DSM in section III-D. This modulated policy, referred to as the transferred policy, is illustrated in Fig. 1, where it also avoids these cups despite the change of goal states. It is worth noting that the transferred constraint can also be utilized by other policy learning/optimization methods such as constrained RL. We adopted DSM in this section only due to its straightforward and rapid modulation process.
@@ V Conclusions and Limitations
This paper proposed the Positive-Unlabeled Constraint Learning method to infer an arbitrary (potentially nonlinear) constraint function from demonstration. The proposed method treats the demonstration as positive data and the higher-reward-winning policy as unlabeled data. It first identifies the reliable infeasible data, and trains a feasibility classifier as constraint function. The benefits of the proposed method were demonstrated learn a 2D or 3D position constraint, using either a constrained RL policy or a DSM policy. It managed to recover and transfer a continuous nonlinear constraint and outperformed other baseline methods in terms of accuracy and safety.
Although PUCL is effective, it only learns a decent but still inaccurate constraint given limited demonstrations and true constraint might be violated in some cases. In the future, we will explore integrating active learning technique to further refine the learned constraint model.

### v2

@@ IV-B Comparison of Policy Representation and Learning Methods
As introduced in section III-D, we studied two policy representation and learning approaches, namely constrained RL and DSM. We test the two variants with PUCL as the constraint learning method in the 3D reaching task. The two frameworks are named CRL-PUCL and DSM-PUCL, respectively, and the results of comparison are presented in Table II. 666 The recall rate is computed as (True infeasible)/(True infeasible + False feasible), and precision is (True infeasible)/(True infeasible + False infeasible) Training was conducted on a PC with a 12th Intel i9-12900K × 24 and an NVIDIA GeForce RTX 3070. All results are the average of 15 independent runs.
@@
DSM-PUCL and CRL-PUCL achieve similar classification performance in IoU, recall rate and precision, and DSM-PUCL generally exhibits lower variance. For the learned policy, while both methods have an unsafe rate below $1\%$, the rate of DSM-PUCL is slightly lower than CRL-PUCL. This is because DSM-PUCL modulates the policy in a principled way and features a safety guarantee [19]. DSM-PUCL also requires less training time than CRL-PUCL. Based on these comparisons, we conclude that both methods are effective for learning constraints. For robotic tasks where the control variable is the velocity/acceleration of each state dimension, DSM may be more favourable than CRL in terms of training time and policy performance, while CRL is more suitable for more general task with more complex environment dynamics and rewards.
TABLE II: Comparison of two policy representation and learning methods
@@
IV-C Constraint Transfer to Variant of the Same Task
One remarkable advantage of learning constraint is that the learned constraint can be transferred to relevant task with the same constraint but potentially different goals and rewards. To demonstrate this, we consider the task of robot avoiding four cups forming two sets of complex shapes (see Fig. 1). We collected demonstrations using DSM and applied DSM-PUCL to learn this constraint. Then the learned constraint network is transferred to a variant task with a set of shifted goal states and starting points. The new nominal policy heading to new target is directly modulated with the transferred constraint with DSM. The modulated policy also avoids these cups despite the change of task configuration.
@@ V Conclusions and Limitations
This paper proposed a two-step Positive-Unlabeled Constraint Learning method to infer an arbitrary, potentially nonlinear, constraint function from demonstration. The proposed method treats the demonstration as positive data and the higher-reward policy as unlabeled data. It uncovers the infeasible states by first identifying reliable infeasible data using a distance metric, and training a feasibility classifier with the identified data. The benefits of the proposed method were demonstrated by learning a 2D and 3D position constraint, a 3D velocity constraint and a constraint in high-dimensional environment, using either a constrained RL policy or a DSM policy. It managed to recover the constraints and outperformed other baseline methods in terms of accuracy and safety.
Although the proposed method is effective, its performance depends on the choice of an appropriate distance threshold and the specific distance metric. In the future, we aim to address this limitation by exploring methods to jointly learn a distance representation that can automatically adapt to various tasks.